# JointDiff: Bridging Continuous and Discrete in Multi-Agent Trajectory Generation

### v1

@@ Abstract
Generative models often treat continuous data and discrete events as separate processes, creating a gap in modeling complex systems where they interact synchronously. To bridge this gap, we introduce JointDiff, a novel diffusion framework designed to unify these two processes by simultaneously generating continuous spatio-temporal data and synchronous discrete events. We demonstrate its efficacy in the sports domain by simultaneously modeling multi-agent trajectories and key possession events. This joint modeling is validated with non-controllable generation and two novel controllable generation scenarios: weak-possessor-guidance, which offers flexible semantic control over game dynamics through a simple list of intended ball possessors, and text-guidance, which enables fine-grained, language-driven generation. To enable the conditioning with these guidance signals, we introduce CrossGuid, an effective conditioning operation for multi-agent domains. We also share a new unified sports benchmark enhanced with textual descriptions for soccer and football datasets. JointDiff achieves state-of-the-art performance, demonstrating that joint modeling is crucial for building realistic and controllable generative models for interactive systems.
@@ 1 Introduction
Modeling the dynamics of multi-agent systems is fundamentally challenging when continuous motion is tightly coupled with discrete, state-altering events. This interplay is critical in domains like autonomous driving and robotics, but finds a particularly rich and demanding testbed in team sports. Here, the continuous trajectories of players are synchronously intertwined with discrete events like passes and possessions. Generating realistic sports gameplay therefore requires a model that can jointly represent these two modalities.
@@
However, existing generative models often fall short by treating these components in isolation. This can lead to physically implausible generations, such as unrealistic passes or flawed ball-possessor interactions (Lee et al., 2024; Capellera et al., 2025). While deterministic models have started to incorporate events (Kim et al., 2023; Capellera et al., 2024), a comprehensive generative framework is missing. This deficiency is compounded by evaluation protocols that rely on individual-level metrics like minimum ADE/FDE (Alahi et al., 2016), which were inherited from pedestrian forecasting and fail to capture scene-level coherence (Casas et al., 2020; Girgis et al., 2021; Weng et al., 2023; Capellera et al., 2025), crucial to team sports.
@@
To address this gap, we turn to the expressive power of diffusion models. While continuous diffusion (Ho et al., 2020) has excelled at generating high-fidelity data like trajectories (Mao et al., 2023; Jiang et al., 2023; Gu et al., 2022; Rempe et al., 2023; Bae et al., 2024; Li et al., 2023; Yang et al., 2024; Capellera et al., 2025), discrete diffusion (Hoogeboom et al., 2021; Austin et al., 2021) has concurrently emerged as a potent, non-autoregressive alternative to large language models (LLMs) for structured sequence generation (Lou et al., 2023). Nascent work has begun to unify these modalities for static tasks such as layout design (Levi et al., 2023) and visual-language modeling (Li et al., 2025). Our key insight is to unify these two paradigms for the temporal-evolving complex systems. We introduce JointDiff, a novel framework that, to the best of our knowledge, is the first to apply joint continuous-discrete diffusion to simultaneously generate spatio-temporal continuous data (trajectories) alongside its corresponding synchronous temporal discrete events (possession events).
@@
Beyond realism, a truly useful generative model also needs to be controllable. While diffusion-guidance methods (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) have been used in motion synthesis to satisfy pedestrian goals or constraints (Jiang et al., 2023; Rempe et al., 2023), semantic control through discrete events in multi-agent systems remains unexplored. Our joint framework directly enables such control using the classifier-free guidance (CFG) (Ho & Salimans, 2022). We introduce weak-possessor-guidance (WPG), a novel conditioning method that allows users to steer gameplay by simply providing an ordered list of intended ball possessors, without rigid timing constraints. We further extend controllability to natural language via text-guided generation, facilitated by a new, curated benchmark of text descriptions for soccer and football datasets. Our approach is illustrated in Fig. 1. In summary, our principal contributions are: 1) A novel joint continuous-discrete diffusion framework that simultaneously generates multi-agent trajectories and synchronous possession events, leading to more realistic and coherent scenes; 2) A thorough evaluation of our joint model on both non-controllable generation and two new controllable tasks: weak-possessor-guidance and text-guidance. For the controllable tasks we introduce CrossGuid, an effective architectural block, to inject the conditioning signal into the denoiser; 3) A unified benchmark for multi-agent modeling in sports, enhanced with new text descriptions for soccer and football datasets. Our method achieves state-of-the-art results on scene-level metrics.
@@
Figure 1: JointDiff. Our model jointly generates continuous trajectories and discrete events, with guidance provided through either weak-possessor information or natural language text. Stars (ˇ) refers to the initial timestep.
@@ 2 Related work
Trajectory Modeling. The evolution of multi-agent trajectory modeling has progressed from Recurrent Neural Networks (RNNs) and Variational RNNs (VRNNs) (Alahi et al., 2016; Felsen et al., 2018; Zhan et al., 2019), to generative models like Generative Adversarial Networks (GANs) (Gupta et al., 2018; Sadeghian et al., 2019; Fang et al., 2020) and Conditional Variational Autoencoders (CVAEs) (Salzmann et al., 2020; Yuan et al., 2021; Lee et al., 2022; Xu et al., 2022). More recently, non-generative transformer-based architectures (Vaswani et al., 2017), sometimes enhanced with visual data (Saadatnejad et al., 2023; Gao et al., 2024), have significantly improved the prediction of multi-modal futures by modeling long-range spatio-temporal dependencies (Girgis et al., 2021; Ngiam et al., 2021; Lee et al., 2024). Building on this progress, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have emerged as the state-of-the-art for generating high-fidelity and diverse trajectories (Mao et al., 2023; Jiang et al., 2023; Gu et al., 2022; Rempe et al., 2023; Bae et al., 2024; Li et al., 2023; Fu et al., 2025; Capellera et al., 2025). This generative power also extends to tasks like trajectory completion, where the recent diffusion model U2Diff (Capellera et al., 2025) has surpassed prior methods based on Graph VRNNs (GVRNNs) (Omidshafiei et al., 2022; Xu et al., 2023), GANs (Liu et al., 2019), and CVAEs (Xu & Fu, 2025). Notably, U2Diff also compete against forecasting-specific architectures despite using an Independent and Identically Distributed (IID) sampling method and without using time window constraints.
@@
Multi-agent Controllability. Diffusion models have recently been augmented with guided sampling to satisfy user-specified constraints or objectives. Existing pedestrian and autonomous driving methods (Rempe et al., 2023; Jiang et al., 2023; Yang et al., 2024), typically focus on controlling individual-level attributes like waypoints, speeds, or physics constraints. Similarly, human motion generation approaches (Karunratanakul et al., 2023) often use waypoints to guide a single human. In contrast, our work focuses on controlling a broader multi-agent system through high-level semantic directives. We adopt the CFG paradigm (Ho & Salimans, 2022), widely used in image and video generation (Rombach et al., 2022; Ho et al., 2022), to bias generation toward a user-specified sequence of possessors or a natural language description. This allows for a comprehensive and intuitive control of the entire scene rather than individual agent behavior.
Joint Continuous-Discrete Diffusion. Jointly modeling continuous and discrete data with diffusion is a growing area, with applications in generative layout design (Levi et al., 2023), event prediction Zeng et al. (2024), CAD sketches (Chereddy & Femiani, 2025), and vision-language modeling (Li et al., 2025). However, this approach remains underexplored for dynamic, spatio-temporal systems. Our work, JointDiff, is the first to introduce this joint modeling paradigm to the domain of time-evolving multi-agent systems, by designing a unified diffusion process that simultaneously generates the continuous trajectories of agents and the discrete sequence of possession events.
@@ 3 Diffusion background
Diffusion models are a class of generative models that learn to reverse a progressive noising process, operating in two stages: a forward diffusion process and a learnable reverse denoising one.
The forward process is a fixed Markov chain that gradually adds noise to a data sample $\bm{\mathrm{X}}_{0}\sim q(\bm{\mathrm{X}}_{0})$. Over $S$ steps, the data is corrupted following a variance schedule, $\{\beta_{s}\in(0,1)\}_{s=1}^{S}$, until $p(\bm{\mathrm{X}}_{S})$ resembles a simple and known noise distribution. This process is defined as: $q(\bm{\mathrm{X}}_{1:S}\mid\bm{\mathrm{X}}_{0})=\prod_{s=1}^{S}q(\bm{\mathrm{X}}_{s}\mid\bm{\mathrm{X}}_{s-1})$. A key property is that we can sample $\bm{\mathrm{X}}_{s}$ at any arbitrary timestep conditioned on the initial data $\bm{\mathrm{X}}_{0}$ in a closed form $q(\bm{\mathrm{X}}_{s}\mid\bm{\mathrm{X}}_{0}$).
The reverse process is a generative model that learns to denoise the data by iteratively reversing the forward steps. Starting with a sample from the known noise distribution, $\bm{\mathrm{X}}_{S}$, a neural network, $p_{\theta}$, learns to approximate the reverse transitions: $p_{\theta}(\bm{\mathrm{X}}_{0:S})=p(\bm{\mathrm{X}}_{S})\prod_{s=1}^{S}p_{\theta}(\bm{\mathrm{X}}_{s-1}\mid\bm{\mathrm{X}}_{s})$.
@@ 4.1 Problem Statement
  • Discrete state: A categorical event at each timestep, such as ball possession, is represented by a one-hot matrix $\bm{\mathrm{E}}\in\{0,1\}^{T\times N}$. Each row $\textbf{e}_{t}$ is a one-hot vector where a value of 1 at index $n$ indicates that agent $n$ is in possession of the ball. States $\textbf{e}_{t}$ where the ball is no possessed (e.g., during a pass or shot) are assigned to the ball’s own category, index $n=0$.
@@ 4.2 Joint Continuous-Discrete Diffusion
where $\lambda$ is a balancing hyperparameter chosen so that both modalities contribute comparably during optimization. Instead of uniform sampling, we use the importance sampling method proposed by Nichol & Dhariwal (2021) to estimate the expectation over the timestep $s$ during training.
@@
Beta Schedule. To align the continuous and discrete sampling processes and improve model accuracy, we employ a hybrid schedule where the total number of discrete steps, denoted as $S^{d}$, is reduced ($S^{d}<S$). Following Levi et al. (2023), we align the discrete steps ($s^{d}$) with the continuous ones $s$ using $s^{d}=\lceil s\cdot(S^{d}/S)\rceil$. We empirically found that a good choice is to match the DDIM skipping step $\zeta$ with the ratio $S/S^{c}$. See an empirical evaluation at Appendix D.3.1.
Controllability. For controllable generation, we utilize the CFG (Ho & Salimans, 2022) training approach. During training, we randomly drop the condition $\mathcal{G}$} with a probability of 25%, which allows the model to learn to denoise both with and without the conditioning information. For non-controllable generation, the model is trained without any conditioning. During inference, we found that we can achieve effective guidance by using a single forward pass with the conditional output.
@@ 4.3 Model Architecture
Our model, JointDiff, builds upon the U2Diff architecture (Capellera et al., 2025), which processes multi-agent trajectories using Social-Temporal Blocks. Each block comprises a Temporal Mamba module (Gu & Dao, 2023) for modeling individual agent dynamics and Social Transformers encoders (Vaswani et al., 2017) for capturing inter-agent interactions. We modify this foundation to create a joint diffusion process capable of controllable generation.
As shown in Fig. 2-left, the model takes as input the noisy state $\bm{\mathrm{X}}_{s}\in{\mathbb{R}}^{T\times N\times 3}$, formed by concatenating the continuous trajectory coordinates $\bm{\mathrm{Y}}_{s}$ and discrete event indicators $\bm{\mathrm{E}}_{s}$ along the feature dimension, and the observed state $\bm{\mathrm{X}}^{\mathrm{co}}\in{\mathbb{R}}^{T\times N\times 3}$ (constructed similarly), along with the binary mask $\textbf{M}\in{\mathbb{R}}^{T\times N}$. It processes these through a series of Social-Temporal Blocks and produces two outputs using:
@@
To enable controllable generation, we introduce the CrossGuid module, which injects an external guidance signal $\mathcal{G}$} into the denoising network. This signal is first encoded into a conditioning tensor $G\in\mathbb{R}^{L\times d}$, where $L$ is the sequence length and $d$ is the feature dimension, both specific to the guidance modality. Refer to Appendix B for architecture details.
Figure 2: Model Architecture. Left: The overall pipeline of our JointDiff model, which takes as input the noisy states $\bm{\mathrm{X}}_{s}$, observed states $\bm{\mathrm{X}}^{\mathrm{co}}$, mask $\mathbf{M}$, and optionally (referred with dashed connections) the encoded guidance signal $G$. Stars (ˇ) refers to the initial timestep, $t=0$. The model outputs the predicted Gaussian noise ${\mathbf{\epsilon}}_{\theta}$ for trajectories and the event probability distribution $\pi_{\theta}$. Right: Detailed view of a Social-Temporal Block featuring our proposed CrossGuid module. The module has two distinct implementations corresponding to different guidance modalities (WPG and Text). The red line (-) in Social-Temporal Block indicates the data flow for non-controllable generation, where CrossGuid is bypassed. An extended diagram is available in Fig. 4.

### v2

@@ Abstract
Generative models often treat continuous data and discrete events as separate processes, creating a gap in modeling complex systems where they interact synchronously. To bridge this gap, we introduce JointDiff, a novel diffusion framework designed to unify these two processes by simultaneously generating continuous spatio-temporal data and synchronous discrete events. We demonstrate its efficacy in the sports domain by simultaneously modeling multi-agent trajectories and key possession events. This joint modeling is validated with non-controllable generation and two novel controllable generation scenarios: weak-possessor-guidance, which offers flexible semantic control over game dynamics through a simple list of intended ball possessors, and text-guidance, which enables fine-grained, language-driven generation. To enable the conditioning with these guidance signals, we introduce CrossGuid, an effective conditioning operation for multi-agent domains. We also share a new unified sports benchmark enhanced with textual descriptions for soccer and football datasets. JointDiff achieves state-of-the-art performance, demonstrating that joint modeling is crucial for building realistic and controllable generative models for interactive systems. Project
@@ 1 Introduction
Modeling the dynamics of multi-agent systems is fundamentally challenging when continuous motion is tightly coupled with discrete, state-altering events. This interplay is critical in domains like autonomous driving and robotics, but finds a particularly rich and demanding testbed in team sports. Here, the continuous trajectories of players are synchronously intertwined with discrete events like passes and possessions. Generating realistic sports gameplay therefore requires a model that can jointly represent these two modalities. However, existing generative models often fall short by treating these components in isolation. This can lead to physically implausible generations, such as unrealistic passes or flawed ball-possessor interactions (Lee et al., 2024; Capellera et al., 2025). While deterministic models have started to incorporate events (Kim et al., 2023; Capellera et al., 2024), a comprehensive generative framework is missing. This deficiency is compounded by evaluation protocols that rely on individual-level metrics like minimum ADE/FDE (Alahi et al., 2016), which were inherited from pedestrian forecasting and fail to capture scene-level coherence (Casas et al., 2020; Girgis et al., 2021; Weng et al., 2023; Capellera et al., 2025), crucial to team sports.
@@
To address this gap, we turn to the expressive power of diffusion models. While continuous diffusion (Ho et al., 2020) has excelled at generating high-fidelity data like trajectories (Mao et al., 2023; Jiang et al., 2023; Gu et al., 2022; Rempe et al., 2023; Bae et al., 2024; Li et al., 2023; Yang et al., 2024; Capellera et al., 2025), discrete diffusion (Hoogeboom et al., 2021; Austin et al., 2021) has concurrently emerged as a potent, non-autoregressive alternative to large language models (LLMs) for structured sequence generation (Lou et al., 2023). Nascent work has begun to unify these modalities for static tasks such as layout design (Levi et al., 2023) and visual-language modeling (Li et al., 2025). Our key insight is to unify these two paradigms for the temporally evolving complex systems. We introduce JointDiff, a novel framework that, to the best of our knowledge, is the first to apply joint continuous-discrete diffusion to simultaneously generate spatio-temporal continuous data (trajectories) alongside its corresponding synchronous temporal discrete events (possession events).
@@
Beyond realism, a truly useful generative model also needs to be controllable. While diffusion-guidance methods (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) have been used in motion synthesis to satisfy pedestrian goals or constraints (Jiang et al., 2023; Rempe et al., 2023), semantic control through discrete events in multi-agent systems remains unexplored. Our joint framework directly enables such control using the classifier-free guidance (CFG) (Ho and Salimans, 2022). We introduce weak-possessor-guidance (WPG), a novel conditioning method that allows users to steer gameplay by simply providing an ordered list of intended ball possessors, without rigid timing constraints. We further extend controllability to natural language via text-guided generation, facilitated by a new, curated benchmark of text descriptions for soccer and football datasets. Our approach is illustrated in Fig. 1. In summary, our principal contributions are: 1) A novel joint continuous-discrete diffusion framework that simultaneously generates multi-agent trajectories and synchronous discrete events, leading to more realistic and coherent scenes; 2) Enabling high-level semantic controllability in dynamic domains. We introduce two novel controllable tasks (weak-possessor-guidance and text-guidance) and a dedicated CrossGuid module that effectively injects conditioning signals into the structured multi-agent embedding; 3) A unified benchmark for multi-agent modeling in sports, enhanced with new text descriptions for soccer and football datasets. Our method achieves state-of-the-art results on scene-level metrics.
@@
Figure 1: JointDiff. Our model jointly generates continuous trajectories and discrete events, with guidance provided through either weak-possessor information or natural language text. Stars ( ) refer to the initial timestep.
@@ 2 Related work
Trajectory Modeling. The evolution of multi-agent trajectory modeling has progressed from Recurrent Neural Networks (RNNs) and Variational RNNs (VRNNs) (Zheng et al., 2016; Felsen et al., 2018; Zhan et al., 2019; Yeh et al., 2019; Li et al., 2021), to generative models like Generative Adversarial Networks (GANs) (Gupta et al., 2018; Sadeghian et al., 2019; Fang et al., 2020) and Conditional Variational Autoencoders (CVAEs) (Salzmann et al., 2020; Graber and Schwing, 2020; Yuan et al., 2021; Lee et al., 2022; Xu et al., 2022; Zheng et al., 2024). In recent years, non-sampling approaches built on transformers (Vaswani et al., 2017), and in some cases enriched with visual data (Saadatnejad et al., 2023; Gao et al., 2024; Wu et al., 2024), have achieved notable progress in multi-modal future prediction by effectively modeling long-range spatio-temporal dependencies (Alcorn and Nguyen, 2021; Girgis et al., 2021; Ngiam et al., 2021; Lee et al., 2024). Building on this progress, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have emerged as the state-of-the-art for generating high-fidelity and diverse trajectories (Mao et al., 2023; Jiang et al., 2023; Gu et al., 2022; Rempe et al., 2023; Bae et al., 2024; Li et al., 2023; Fu et al., 2025). This generative power also extends to tasks like trajectory completion, where the recent diffusion model U2Diff (Capellera et al., 2025) has surpassed prior methods based on Graph VRNNs (GVRNNs) (Omidshafiei et al., 2022; Xu et al., 2023), GANs (Liu et al., 2019), and CVAEs (Xu and Fu, 2025). Notably, U2Diff also competes against forecasting-specific architectures despite using an Independent and Identically Distributed (IID) sampling method and without using time window constraints.
@@
Multi-agent Controllability. Diffusion models have recently been augmented with guided sampling to satisfy user-specified constraints or objectives. Existing pedestrian and autonomous driving methods (Rempe et al., 2023; Jiang et al., 2023; Yang et al., 2024), typically focus on controlling individual-level attributes like waypoints, speeds, or physics constraints. Similarly, human motion generation approaches (Karunratanakul et al., 2023) and robotics planning methods (Mishra et al., 2023; Fang et al., 2024) often guide a single agent. In contrast, our work focuses on controlling a broader multi-agent system through high-level semantic directives. We adopt the CFG paradigm (Ho and Salimans, 2022), widely used in image and video (Rombach et al., 2022; Ho et al., 2022), to bias generation toward a user-specified sequence of possessors or a natural language description. This allows for a comprehensive control of the entire scene rather than individual agent behavior.
Joint Continuous-Discrete Diffusion. Joint diffusion models for mixed continuous–discrete data are an emerging research direction, with applications in static domains such as layout design (Levi et al., 2023), CAD sketches (Chereddy and Femiani, 2025), and vision–language modeling (Li et al., 2025), where absorbing state diffusion (Austin et al., 2021) is commonly used for discrete variables. In contrast, dynamic domains have been underexplored. Prior work (Zeng et al., 2024) applies the multinomial formulation (Hoogeboom et al., 2021) to temporal point processes, but it’s restricted to single-instance future prediction and relies on sequential, non-simultaneous generation. We extend the multinomial formulation to general controllable dynamic domains, exemplified by the multi-agent completion task, and introduce a unified diffusion framework that simultaneously models continuous trajectories and discrete events. This formulation proves more consistent than absorbing state diffusion in our temporally evolving domain, as it enables continuous refinement of discrete variables throughout the denoising process. Furthermore, we are first to incorporate high-level semantic controllability, such as WPG and text-guidance, for joint continuous–discrete generation in dynamic domains, consistently outperforming the non-joint baseline.
@@ 3 Diffusion background
Diffusion models are a class of generative models that learn to reverse a progressive noising process, operating in two stages: a forward diffusion process and a learnable reverse denoising process. The forward process is a fixed Markov chain that gradually adds noise to a data sample $\bm{\mathrm{X}}_{0}\sim q(\bm{\mathrm{X}}_{0})$. Over $S$ steps, the data is corrupted following a variance schedule, $\{\beta_{s}\in(0,1)\}_{s=1}^{S}$, until $p(\bm{\mathrm{X}}_{S})$ resembles a simple and known noise distribution. This process is defined as: $q(\bm{\mathrm{X}}_{1:S}\mid\bm{\mathrm{X}}_{0})=\prod_{s=1}^{S}q(\bm{\mathrm{X}}_{s}\mid\bm{\mathrm{X}}_{s-1})$. A key property is that we can sample $\bm{\mathrm{X}}_{s}$ at any arbitrary timestep conditioned on the initial data $\bm{\mathrm{X}}_{0}$ in a closed form $q(\bm{\mathrm{X}}_{s}\mid\bm{\mathrm{X}}_{0}$). The reverse process is a generative model that learns to denoise the data by iteratively reversing the forward steps. Starting with a sample from the known noise distribution, $\bm{\mathrm{X}}_{S}$, a neural network, $p_{\theta}$, learns to approximate the reverse transitions: $p_{\theta}(\bm{\mathrm{X}}_{0:S})=p(\bm{\mathrm{X}}_{S})\prod_{s=1}^{S}p_{\theta}(\bm{\mathrm{X}}_{s-1}\mid\bm{\mathrm{X}}_{s})$.
@@ 4.1 Problem Statement
  • Discrete state: A categorical event at each timestep, such as ball possession, is represented by a one-hot matrix $\bm{\mathrm{E}}\in\{0,1\}^{T\times N}$. Each row $\textbf{e}_{t}$ is a one-hot vector where a value of 1 at index $n$ indicates that agent $n$ is in possession of the ball. States $\textbf{e}_{t}$ where the ball is not possessed (e.g., during a pass or shot) are assigned to the ball’s own category, index $n=0$.
@@ 4.2 Joint Continuous-Discrete Diffusion
where $\lambda$ is a balancing hyperparameter chosen so that both modalities contribute comparably during optimization. Instead of uniform sampling, we use the importance sampling method proposed by Nichol and Dhariwal (2021) to estimate the expectation over the timestep $s$ during training.
@@
Beta Schedule. To align the continuous and discrete sampling processes and improve model accuracy, we employ a hybrid schedule where the total number of discrete steps, denoted as $S^{d}$, is reduced ($S^{d}<S$). Following Levi et al. (2023), we align the discrete steps ($s^{d}$) with the continuous ones $s$ using $s^{d}=\lceil s\cdot(S^{d}/S)\rceil$. We empirically found that a good choice is to match the DDIM skipping step $\zeta$ with the ratio $S/S^{d}$. See an empirical evaluation at Appendix D.3.1.
Controllability. For controllable generation, we utilize the CFG (Ho and Salimans, 2022) training approach. During training, we randomly drop the condition $\mathcal{G}$} with a probability of 25%, which allows the model to learn to denoise both with and without the conditioning information. For non-controllable generation, the model is trained without any conditioning. During inference, we found that we can achieve effective guidance by using a single forward pass with the conditional output. See an empirical evaluation at Appendix D.3.3.
@@ 4.3 Model Architecture
Our model, JointDiff, builds upon the U2Diff architecture (Capellera et al., 2025), which processes multi-agent trajectories using Social-Temporal Blocks. Each block comprises a Temporal Mamba module (Gu and Dao, 2023) for modeling individual agent dynamics and Social Transformers encoders (Vaswani et al., 2017) for capturing inter-agent interactions. We modify this foundation to create a joint diffusion process capable of controllable generation.
As shown in Fig. 2-left, the model takes as input the noisy state $\bm{\mathrm{X}}_{s}\in{\mathbb{R}}^{T\times N\times 3}$, formed by concatenating the continuous trajectory coordinates $\bm{\mathrm{Y}}_{s}$ and discrete event indicators $\bm{\mathrm{E}}_{s}$ along the feature dimension, and the observed state $\bm{\mathrm{X}}^{\mathrm{co}}\in{\mathbb{R}}^{T\times N\times 3}$ (constructed similarly), along with the binary mask $\textbf{M}\in{\mathbb{R}}^{T\times N}$. It processes these through two Social-Temporal Blocks and produces two outputs using:
@@
To enable controllable generation, we introduce the CrossGuid module, which injects an external guidance signal $\mathcal{G}$} into the denoising network. This signal is first encoded into a conditioning tensor $G\in\mathbb{R}^{L\times d}$, where $L$ is the sequence length and $d$ is the feature dimension, both specific to the guidance modality. During training, conditioning dropout ($\mathcal{G}$$\emptyset$) is performed by setting $G$ to a zero tensor. Refer to Appendix B for architecture details.
Figure 2: Model Architecture. Left: The overall pipeline of our JointDiff model, which takes as input the noisy states $\bm{\mathrm{X}}_{s}$, observed states $\bm{\mathrm{X}}^{\mathrm{co}}$, mask $\mathbf{M}$, and optionally (referred with dashed connections) the encoded guidance signal $G$. Stars ( ) refer to the initial timestep, $t=0$. The model processes these inputs through two Social-Temporal Blocks and outputs the predicted Gaussian noise ${\mathbf{\epsilon}}_{\theta}$ for trajectories and the event probability distribution $\pi_{\theta}$. Right: Detailed view of a Social-Temporal Block featuring our proposed CrossGuid module. The module has two distinct implementations corresponding to different guidance modalities (WPG and Text). The red line (-) in Social-Temporal Block indicates the data flow for non-controllable generation, where CrossGuid is bypassed. An extended diagram is available in Fig. 5.